# A Morse-Bott unification of the Grassmannians of a symplectic vector space

### v1

@@ 1 Introduction
At the linear level, this relationship is captured by the Kähler angle $\theta_{W}$ of a $2$-plane $W$ (see [13] and references therein), which measures the deviation of $W$ from being a complex line. It is well known that $\theta_{W}=0$ when $W$ is complex and $\theta_{W}=\pi/2$ when $W$ is isotropic. While Harvey-Lawson [5] gives a multiplicative generalization, we present a distinct, concrete geometric model of this interaction using Morse-Bott theory. Our Morse-Bott function organizes all linear subspaces of a symplectic vector space according to their compatibility with $J$, and is an additive, multidimensional generalization of the Kähler angle associated to it (Proposition 3.3).
@@
This function measures the failure of the subspace to be $J$-invariant. We also interpret this quantity dynamically (Lemma 3.1) as the kinetic energy of the fundamental vector field $X_{J}$ generated by the $U(1)$-action of $J$ on the Grassmannian:
@@
Analytically, $f$ mimics the norm-squared of a moment map in the sense of [6] (for other Morse functions on Grassmannians see [4], [15], [10]). Although the real Grassmannian is not symplectic in general (for dimensional reasons) and thus admits no classical moment map, $f$ also induces a manifold decomposition invariant under its symmetries. Remarkably, while $f$ is defined via the complex structure J, its stable manifolds recover the orbit decomposition of the real Grassmannian $\operatorname{Gr}(k,V)$ under the linear symplectomorphism group $\operatorname{Sp}(V)$.
These orbits are labeled by the invariant indices $\vec{n}=(n_{0},n_{+},n_{-})$, corresponding to the types of linear subspaces: the dimensions of the isotropic ($n_{0}$) and reduced ($2n_{+}$) parts of $W$, and the reduced part of its complement ($2n_{-}$). We view these orbits as Grassmannian manifolds $\operatorname{Gr}(\vec{n};V)$, generalizing the Lagrangian ($n_{0}=n$), isotropic ($n_{+}=0$), coisotropic ($n_{-}=0$), and symplectic ($n_{0}=0$) Grassmannians.
@@
This result unifies several topological observations scattered throughout the symplectic literature. Arnold [3] described the Lagrangian Grassmannian as $\operatorname{U}(n)/\operatorname{O}(n)$; Oh-Park [11] described the coisotropic case; and Lee-Leung [8] and Ajayi-Banyaga [1] analyzed the symplectic case. Our corollary resolves the homotopy type for the ”mixed” cases (neither isotropic, coisotropic, nor symplectic) and constructs all deformation retracts simultaneously via a single Morse-theoretic framework.
We note that this retraction is not easily accessible via standard symmetric space theory (e.g., [9], Theorem IV.3.5). The stabilizer subgroups $\operatorname{Sp}_{W}(V)$ are nonreductive, as they generally contain nilpotent subgroups (Proposition 2.6), so they cannot be realized as the fixed points of an involution. While Takeuchi [14] used negative gradient flows to construct retractions in Hermitian symmetric spaces, that construction was restricted to open orbit strata. Our method applies to all strata uniformly.
@@ 2.1 Types of linear subspaces
$W^{\omega}=\{v\in V\mid\omega(v,w)=0,\forall w\in W\}.$ (4)
@@
We classify subspaces by their type, a triple of nonnegative integers $\vec{n}=(n_{0},n_{+},n_{-})$ defined by:
@@ 2.2 Transitivity of the linear symplectomorphism group
Theorem 2.2 (Linear relative Darboux theorem ([2])).
@@ 2.3 Stabilizer Subgroups and Splittings
We use the matrix representation in a Darboux basis from Theorem 2.2. In this Darboux basis, an element $g\in H(W)$ takes the block form
@@
However, for general isotropic subspaces ($n_{+}>0$ or $n_{-}>0$), the group $H(W)$ is non-abelian and 2-step nilpotent. It is isomorphic to a generalized Heisenberg group [16]. Specifically, the center of $H(W)$ corresponds to the symmetric forms on $W_{0}$, while the quotient by the center acts as a translation group on the symplectic components. The commutator bracket in the Lie algebra $\mathfrak{h}(W)$ reproduces the symplectic form pairing, reflecting the canonical central extension structure defining Heisenberg groups.
@@ 2.5 $J$-compatible subspaces
Conclusion: Combining these results, the stabilizer is isomorphic to the product of these groups:
@@ 3.1 Differential geometry of the real Grassmannian
We fix an identification $(V,\omega,J)\cong\mathbb{R}^{2n}$. Denote the inner product $g(u,v):=\omega(Ju,v)$ and norm $|u|^{2}=(u,u)$. We identify the real Grassmannian $\operatorname{Gr}(k;V)$ with its image under the embedding $\operatorname{Gr}(k;V)\hookrightarrow\operatorname{Mat}_{2n\times 2n}(\mathbb{R})$ mapping a linear subspace $W$ to its orthogonal projection matrix $P_{W}$. The ambient space $\operatorname{Mat}_{2n\times 2n}(\mathbb{R})$ has an $\operatorname{O}(2n)$-invariant trace inner product $\langle A,B\rangle=\operatorname{Tr}(AB^{T})$, and $\operatorname{Gr}(k;V)$ becomes a Riemannian manifold endowed with the restricted trace inner product. We denote the norm $\lVert A\rVert^{2}:=\langle A,A\rangle$ and the matrix commutator by $A,B]=\operatorname{ad}_{A}(B)$.
@@
Thus we will identify the tangent space at $P$ as the +1+1-eigenspace of the matrix anticommutator $\{P,\cdot\}$:
$T_{P}\operatorname{Gr}(k;V)\cong\{X\in\Sigma:\{P,X\}=X\}\subseteq\operatorname{Mat}_{2n\times 2n}(\mathbb{R}).$
@@
    Let $P(t)$ be a curve in $\operatorname{Gr}(k;V)$ with $P(0)=P$ and $\dot{P}(0)=Y$. Let $X(t)$ be a section of the ambient tangent bundle $T\Sigma|_{\operatorname{Gr}(k;V)}$ along $P(t)$. The shape operator is $A_{X}(Y)=-\operatorname{pr}_{P}(\dot{X}(0))$.
@@
    Since $S$ commutes with $[Y,X]$, $P=\frac{1}{2}(S+\operatorname{Id})$ also commutes with $[Y,X]$. Thus, $A_{X}(Y)=[P,[Y,X]]=0$.
@@ 3.2 The energy function and the Kähler angle
$f(P):=\frac{1}{2}\lVert X_{J}(P)\rVert^{2}=\frac{1}{2}\operatorname{Tr}\left([P,J][P,J]^{T}\right)=\frac{1}{2}\operatorname{Tr}([P,J]^{2})$
@@
    Recall On each block $V_{j}^{\theta}$, let $W_{j}^{\theta}=W\cap V_{j}^{\theta}$. Since $\dim W_{j}^{\theta}=2$, we have $\operatorname{Tr}(P|_{V_{j}^{\theta}})=2$. The operator $PJPJ$ restricted to $W_{j}^{\theta}$ acts as $(\cos\theta_{j})^{2}\operatorname{Id}$. Thus,
@@ 3.4 The Morse-Bott property
$\operatorname{ad}_{P}(\mathfrak{u}(V,J))=T_{P}\operatorname{Gr}_{J}(\vec{n};V)=\ker H(f)_{P},$
@@ 3.5 Gradient flows
Since $\operatorname{Gr}(k;V)$ is compact, the negative gradient flow $\{\phi_{t}\}_{t\in\mathbb{R}}$ of $f$ is complete and $\lim_{t\to\infty}\phi_{t}(P)$ exists. Denote by $P(t):=\phi_{t}(P)$ and $P_{\infty}:=\lim_{t\to\infty}\phi_{t}(P)$.
@@
where $\phi_{t}^{j}$ is the negative gradient flow of $f_{j}$ on $\operatorname{Gr}(\dim W_{j};V_{j})$.
@@ 3.6 Type preservation
Multiplying by $P^{\perp}=(I-P)$ on the left annihilates the second term ($P^{\perp}P=0$), yielding:
@@
$X_{\xi}(P)=\dot{P}=\frac{1}{2}\left[P,\left[P,\xi+\xi^{T}\right]\right]-\frac{1}{2}\left[P,\xi-\xi^{T}\right]=-\operatorname{ad}_{P}^{\bullet}(\xi).$
The kernel of $\xi\mapsto X_{\xi}(P)$ is the stabilizer algebra $\mathfrak{sp}_{W}(V)$ (generators fixing $W$). The map is surjective onto the tangent space of the orbit $T_{W}(\operatorname{Sp}(V)\cdot W)$. Thus the sequence is exact. ∎
@@
$X_{Z}(P)=-\operatorname{ad}^{\bullet}_{P}(Z)=-(-\operatorname{ad}_{P}^{2}(Z))=[P,[P,Z]].$
@@
For any point $P$ in $\operatorname{Gr}(k;V)$,
$n_{0}(P)=n_{0}(P_{\infty}).$
@@
Let $P\in\operatorname{Gr}(\vec{n};V)$. By Lemma 3.13, the flow trajectory lies entirely within the $\operatorname{Sp}(V)$-orbit of $P$. Thus, the limit $P_{\infty}$ must lie in the closure of this orbit:
@@
By Proposition 3.7, we have $n_{0}(P_{\infty})\geq n_{0}(P)$.
Since $f(P(t))$ is monotonically decreasing
$f(P_{\infty})\leq f(P(t))$
for all $t$. By Proposition 3.7, the left hand side is $f(P_{\infty})=n_{0}(P_{\infty})$.
By Proposition 3.3 and Lemma 3.11, the right hand bound sharpens to $n_{0}(P)$ as $t\to\infty$. (In each $V_{j}^{\theta}$ of Proposition 2.13, Lemma 3.11 gives $f_{j}(W_{j}(t))\to 0$; summing over $j$ yields $f(P(t))\to n_{0}(P)$.) So $n_{0}(P_{\infty})\leq n_{0}(P)$. ∎

### v2

@@ 1 Introduction
At the linear level, this relationship is captured by the Kähler angle $\theta_{W}$ of a $2$-plane $W$ (Definition 2.11 or see [22] and references therein), which measures the deviation of $W$ from being a complex line. It is well known that $\theta_{W}=0$ when $W$ is complex and $\theta_{W}=\pi/2$ when $W$ is isotropic. While Harvey-Lawson [11] and Donaldson [7] work with a multiplicative generalization, we present an additive generalization to higher dimensions (Proposition 3.3), together with a concrete geometric model via Morse-Bott theory [6, 4]. Our Morse-Bott function assigns a number to all linear subspaces of a symplectic vector space according to their compatibility with $J$.
@@
This function measures the failure of the subspace to be $J$-invariant. We also interpret this quantity dynamically (Lemma 3.1) as the kinetic energy of the fundamental vector field $X_{J}$ generated by the $\operatorname{U}(1)$-action of $J$ on the Grassmannian:
@@
Analytically, $f$ mimics the norm-squared of a moment map in the sense of [13] (for linear Morse functions on Grassmannians see [10, 24, 8, 9, 5, 18]). Although the real Grassmannian is not symplectic in general (for dimensional reasons) and thus admits no classical moment map, $f$ also induces a manifold decomposition invariant under its symmetries. Remarkably, while $f$ is defined via the complex structure $J$, its stable manifolds recover the orbit decomposition of the real Grassmannian $\operatorname{Gr}(k;V)$ under the linear symplectomorphism group $\operatorname{Sp}(V)$.
These orbits are labeled by the invariant indices (Section 2.1) $\vec{n}=(n_{0},n_{+},n_{-})$, which we term the type of the subspace, in analogy with Harald Hess’s classification of polarizations (Definition 3.1 of [12]). These indices correspond to the dimensions of the isotropic ($n_{0}$) and reduced ($2n_{+}$) parts of $W$, and the reduced part of its complement ($2n_{-}$). We view these orbits as Grassmannian manifolds $\operatorname{Gr}(\vec{n};V)$, generalizing the Lagrangian ($n_{0}=n$), isotropic ($n_{+}=0$), coisotropic ($n_{-}=0$), and symplectic ($n_{0}=0$) Grassmannians.
@@
This result unifies several topological observations scattered throughout the symplectic literature. Arnold [3] described the Lagrangian Grassmannian as $\operatorname{U}(n)/\operatorname{O}(n)$; Oh-Park [19] described the coisotropic case; and Lee-Leung [15] and Ajayi-Banyaga [1] analyzed the symplectic case. Our corollary resolves the homotopy type for the “mixed” cases (neither isotropic, coisotropic, nor symplectic) and constructs all deformation retracts simultaneously via a single Morse-theoretic framework.
We note that this retraction is not easily accessible via standard methods (e.g., [17, Theorem IV. 3.5]). The stabilizer subgroups $\operatorname{Sp}_{W}(V)$ are nonreductive, as they generally contain nilpotent subgroups (Proposition 2.6), so they cannot be realized as the fixed points of an involution. This nonreductivity limits the use of the combinatorics of root systems and GIT quotients. While Takeuchi [23] used negative gradient flows to construct retractions in Hermitian symmetric spaces, that construction was restricted to open orbit strata (where the stabilizer is reductive).
@@ 2.1 Types of linear subspaces
$W^{\omega}=\{v\in V\mid\omega(v,w)=0,\text{ for all }w\in W\}.$ (4)
@@
We adapt the notation of Harald Hess [12, Definition 3.1], and classify subspaces by their type, a triple of nonnegative integers $\vec{n}=(n_{0},n_{+},n_{-})$ defined by:
@@ 2.2 Transitivity of the linear symplectomorphism group
Theorem 2.2 (Linear relative Darboux theorem ([2, Section 1.2])).
@@ 2.3 Stabilizer Subgroups and Splittings
We take a Darboux basis from Theorem 2.2, and reorder the blocks as
$\{e^{0},e^{+},f^{+},e^{-},f^{-},f^{0}\}.$
In this ordered basis, an element $g\in H(W)$ takes the block form
@@
However, for general isotropic subspaces ($n_{+}>0$ or $n_{-}>0$), the group $H(W)$ is non-abelian and 2-step nilpotent. It is isomorphic to a generalized Heisenberg group [25]. Specifically, the center of $H(W)$ corresponds to the symmetric forms on $W_{0}$, while the quotient by the center acts by shearing the symplectic components $W_{\pm}$ along $W_{0}$.
@@ 2.5 $J$-compatible subspaces
Combining these results, the stabilizer is isomorphic to the product of these groups:
@@ 3.1 Differential geometry of the real Grassmannian
We fix an identification $(V,\omega,J)\cong\mathbb{R}^{2n}$. Denote the inner product $g(u,v):=\omega(Ju,v)$ and norm $|u|^{2}=g(u,u)$. We identify the real Grassmannian $\operatorname{Gr}(k;V)$ with its image under the embedding $\operatorname{Gr}(k;V)\hookrightarrow\operatorname{Mat}_{2n\times 2n}(\mathbb{R})$ mapping a linear subspace $W$ to its orthogonal projection matrix $P_{W}$. The ambient space $\operatorname{Mat}_{2n\times 2n}(\mathbb{R})$ has an $\operatorname{O}(2n)$-invariant trace inner product $\langle A,B\rangle:=\operatorname{Tr}(AB^{T})$, and $\operatorname{Gr}(k;V)$ becomes a Riemannian manifold endowed with the restricted trace inner product. We denote the norm $\lVert A\rVert^{2}:=\langle A,A\rangle$ and the matrix commutator by $A,B]=\operatorname{ad}_{A}(B)$.
@@
Thus we will identify the tangent space at $P$ as the kernel of the matrix anticommutator with the reflection matrix $\{2P-\operatorname{Id},\cdot\}$:
$T_{P}\operatorname{Gr}(k;V)\cong\{X\in\Sigma:\{2P-\operatorname{Id},X\}=0\}.$
@@
    Let $P(t)$ be a curve in $\operatorname{Gr}(k;V)$ with $P(0)=P$ and $\dot{P}(0)=Y$. Let $X(t)$ be a section of the ambient tangent bundle $T\Sigma|_{\operatorname{Gr}(k;V)}$ along $P(t)$. The shape operator is $\mathcal{A}_{X}(Y)=-\operatorname{pr}_{P}(\dot{X}(0))$.
@@
    Since $S$ commutes with $[Y,X]$, $P=\frac{1}{2}(S+\operatorname{Id})$ also commutes with $[Y,X]$. Thus, $\mathcal{A}_{X}(Y)=[P,[Y,X]]=0$.
@@ 3.2 The energy function and the Kähler angle
$f(P):=\frac{1}{2}\lVert(X_{J})_{P}\rVert^{2}=\frac{1}{2}\operatorname{Tr}\left([P,J][P,J]^{T}\right)=\frac{1}{2}\operatorname{Tr}([P,J]^{2})$
@@
    On each block $V_{j}^{\theta}$, let $W_{j}^{\theta}=W\cap V_{j}^{\theta}$. Since $\dim W_{j}^{\theta}=2$, we have $\operatorname{Tr}(P|_{V_{j}^{\theta}})=2$. Recall from the proof of Theorem 2.12 that the Kähler angle $\theta_{j}$ is defined via the operator $S=-(PJ|_{W})^{2}$, whose eigenvalues on $W_{j}^{\theta}$ are $(\cos\theta_{j})^{2}$. Since $PJPJ|_{W_{j}^{\theta}}=-S$, this operator acts as $(\cos\theta_{j})^{2}\operatorname{Id}$. Thus,
@@ 3.4 The Morse-Bott property
$\operatorname{ad}_{P}(\mathfrak{u}(V,J))=T_{P}\operatorname{Gr}_{J}(\vec{n};V)=\ker\operatorname{Hess}(f)_{P},$
@@ 3.5 Gradient flows
Since $\operatorname{Gr}(k;V)$ is compact, the negative gradient flow $\phi_{t}$ exists for all $t\in\mathbb{R}$. Moreover, since $f$ is real analytic, the Łojasiewicz gradient inequality ensures that these flow lines have finite length and each flow line converges to a unique critical point [16], [20, Section IV.9].
We denote the trajectory by $P(t):=\phi_{t}(P)$ and its limit by $P_{\infty}:=\lim_{t\to\infty}P(t)$.
@@
where $\phi_{t}^{j}$ is the negative gradient flow of $f_{j}$ on $\operatorname{Gr}(2;V_{j}^{\theta})$.
@@ 3.6 Type preservation
Multiplying by $P^{\perp}=\operatorname{Id}-P$ on the left annihilates the second term ($P^{\perp}P=0$), yielding:
@@
$(X_{\xi})_{P}=\dot{P}=\frac{1}{2}\left[P,\left[P,\xi+\xi^{T}\right]\right]-\frac{1}{2}\left[P,\xi-\xi^{T}\right]=-\operatorname{ad}_{P}^{\bullet}(\xi).$
The kernel of $\xi\mapsto(X_{\xi})_{P}$ is the stabilizer algebra $\mathfrak{sp}_{W}(V)$ (generators fixing $W$). The map is surjective onto the tangent space of the orbit $T_{W}(\operatorname{Sp}(V)\cdot W)$. Thus the sequence is exact. ∎
@@
$(X_{Z})_{P}=-\operatorname{ad}^{\bullet}_{P}(Z)=-(-\operatorname{ad}_{P}^{2}(Z))=[P,[P,Z]].$
@@
Let $\{\phi_{t}\}_{t\in\mathbb{R}}$ be the negative gradient flow for the energy function, and for any point $P$ in $\operatorname{Gr}(k;V)$, let $P(t):=\phi_{t}(P)$ and $P_{\infty}:=\lim_{t\to\infty}P(t)$. Then
$n_{0}(P)=n_{0}(P(t))=n_{0}(P_{\infty}).$
@@
Let $P\in\operatorname{Gr}(\vec{n};V)$. By Lemma 3.13, the gradient vector field is tangent to the symplectic orbit $\operatorname{Sp}(V)\cdot P$. Therefore, for all finite $t$, the flow $P(t)$ remains in the same orbit, implying $n_{0}(P(t))=n_{0}(P)$.
Moreover, the limit $P_{\infty}$ must lie in the closure of this orbit (Proposition 2.3):
@@
Thus $n_{0}(P_{\infty})\geq n_{0}(P)$.
Since $f(P(t))$ is monotonically decreasing $f(P_{\infty})\leq f(P(t))$ for all $t$. By Proposition 3.7, the left hand side is $f(P_{\infty})=n_{0}(P_{\infty})$.
By Proposition 3.3 and Lemma 3.11, the right hand bound sharpens to $n_{0}(P)$ as $t\to\infty$. (In each four dimensional symplectic vector space $V_{j}^{\theta}$ of Proposition 2.13, Lemma 3.11 shows the energy decays as $f_{j}(W_{j}(t))\to 0$; summing over $j$ yields $f(P(t))\to n_{0}(P)$.) So $n_{0}(P_{\infty})\leq n_{0}(P)$. ∎